OpenStudy (anonymous):

What is y=2x squared-4x-1... when x equals -2, -1, 0, 1, 2, 3.

OpenStudy (angela210793):

y=2x^2-4x-1 whn x=-2 2*(-2)^2-4*(-2)-1=8+8-1=15 put -1, 0, 1, 2, 3 instead of x,multiply and add